2. Suppose the market supply function for natural gas is P = 10 + 2Q and the market demand function of natural gas is P = 70 - Q, where P is the price of the natural gas per cubic feet and Q is the quantity of natural gas bought and sold. 1) What are the equilibrium price and quantity of natural gas in a competitive market? 2) Compute the consumer surplus and producer surplus. Assume the government imposes a price ceiling at P = $40. 3) Find the consumer surplus and producer surplus associated with the resulting allocation. 4) Find total benefit and total cost in the allocation.
Answer #1

1.2.1) Demand: P: 70-a Suppry: P: 10+2 Q. Equilibrium: to. Q = 10+2Q or, 30 = 60 ..0 =20} : :. p = 501 yox . X 15 20 Jo . 2)